# Using Large Language Models for an ISO37101 standard assessment mapping for sustainable communities

### v1

@@ 2.1 Sustainable urban development and assessment frameworks
The concept of sustainable urban development has evolved considerably since its beginnings in the late 20th century. Subsequently, frameworks such as the United Nations Sustainable Development Goals (SDGs) have set global targets, with SDG 11 specifically addressing sustainable cities and communities. Different initiatives such as CASBEE for Urban Development, LEED for Neighbourhoods and BREEAM Communities [3, 4, 5] provide standardised criteria for assessing sustainable urban development at the neighbourhood level [6, 7], which are helpful in evaluating initiatives to achieve certification [8], but from a strategic perspective, these frameworks often focus primarily on specific (social or environmental) aspects and may overlook broader social and economic dimensions of sustainability
@@ 2.2 ISO 37101: Structure, purpose, and applications
The series containing ISO 37101 [15] , which was introduced in 2016, represents a more holistic approach to assessing urban sustainability, including mapping to a set of urban indicators [16], specified for example in ISO7310 and operationalized in cities indicators datasets [17, 18]. Its structure is based on six purposes (attractiveness, preservation and improvement of the environment, resilience, responsible resource use, social cohesion, and well-being) and twelve issues (from governance to biodiversity). provides a comprehensive matrix for assessing urban initiatives, as well as specific indicators [19]. This approach was recently explored by the French [20, 21]. Thanks to the flexibility of the standard, it can be adapted to different scales, from individual projects to city-wide strategies. However, the relative complexity of the framework can make widespread adoption difficult, especially in resource-constrained urban environments
@@ 3 Methodology
We selected gpt-3.5-turbo as the primary LLM, with additional tests using other local models (llama2, mistral, …). The choice of gpt-3.5-turbo was based on its proven ability to understand complex relationships and generate nuanced responses, as well as its ability to use “functions calling” to obtain JSON-structured data. This approach can later be ported to more recent models (in particular the latest OpenAI 4o-mini model). Its large knowledge base, trained on various datasets, makes it well suited for interpreting different urban project descriptions from a general perspective. We also considered factors such as the accessibility of the API, the consistency of the results and the possibilities for fine-tuning. We tried locally-runned models as secondary models to allow to compare the results and qualitatively assess the generalisability of our approach across different LLMs, but rapidly discarded this because of the speed and usability across large datasets.
A critical component of this LLM-based approach the development of a custom prompt that summarises the essence of the ISO 37101 framework. This prompt was iteratively refined to guide the LLM in the use of the 6x12 framework. The prompt includes concise definitions of the six sustainability purposes and twelve issues, instructions for identifying relevant aspects of the input text, Guidelines for assessing the project’s contribution to each purpose-issue intersection, and a structured format for the output of the assessment results, qualified in the functions calling part of the request. Prompts have iteratively been refined through testing with a variety of urban project descriptions to ensure consistency and accuracy of LLM interpretations.
@@
  • Creation of visualisations and reports in Excel format.
4 Case Studies and Results
@@ 4.1.2 Qualitative analysis and control of LLM outputs
Figure 1: Examples of reviews. On the left hand side, a review from an urban project in Grenoble [26]. On the right hand side, review of PROBONO [25] activities.
@@ 4.2.1 Paris
Moreover, based on this dataset, we developed a catalogue of "sustainable ideas" in which we have categorised exemplary projects under each purpose-issue overlap. This catalogue serves as a resource for urban planners and citizens and provides inspiration for future participatory budgeting proposals focused on sustainability goals.
@@ 4.2.2 The PROBONO project
FIGURE FROM PROBONO
@@ 5.2 Comparison with traditional human-driven assessment methods
The advantages of an LLM-based assistant versus fully human reviews are twofold. The first is the significant time saving compared to traditional manual assessments (a few seconds compared to around 30 minutes per project for an overview of the project). This efficiency gain opens up opportunities for more comprehensive and frequent sustainability assessments, allowing communities to assess a much wider range of initiatives and update their assessments more regularly. The second aspect is the consistency of automated assessments, which offsets another limitation of manual methods, namely the potential variability between assessors. This approach provides a standardised interpretation that can be particularly useful for comparing projects in different contexts or time periods. Tangentially, LLM have also enabled the use of virtual citizens [28], with obvious limitations.
Nevertheless, the use of AI in urban planning decision-making processes raises important ethical considerations [29]. It is essential to ensure that LLMs complement, rather than replace, human expertise. Furthermore, while LLMs are excellent at recognising explicit relationships, they occasionally overlook subtle, context-specific factors that can be easily identified by human experts. There is therefore a risk that important qualitative factors or local knowledge that cannot be captured in standardised frameworks will be marginalised.
@@ 5.5 Applicability to solutions identification
This exchange of ideas and solutions has the potential to be powerful as it is based on the overview understanding of city services and sustainability impacts rather than superficial similarities. For example, a community’s need for greater social cohesion around environmental preservation may lead to the discovery of innovative green space initiatives that have fostered community engagement elsewhere. This approach can accelerate the learning process for communities stakeholder, and the adaptation and scaling of proven solutions in different urban contexts. It also encourages a collaborative and globally focussed approach to urban problem solving, where communities can learn from and build on the experiences of others.
@@ 7 Conclusion
Furthermore, the adaptability of our approach in different contexts - from open citizen data in the Paris participatory budget to specific project information in the context of PROBONO activities - demonstrates its versatility and potential for wide-ranging applications in urban planning and governance. This flexibility suggests that the method could be useful for cities of different sizes, cultural contexts and stages of development.

### v2

# Using Large Language Models for a standard assessment mapping for sustainable communities

@@ 2.1 Sustainable urban development and assessment frameworks
The concept of sustainable urban development has evolved considerably since its beginnings in the late 20th century. Subsequently, frameworks such as the United Nations Sustainable Development Goals (SDGs) have set global targets, with SDG 11 specifically addressing sustainable cities and communities. Different initiatives such as CASBEE for Urban Development, LEED for Neighbourhoods, BREEAM Communities [3, 4, 5] and DGNB Urban Districts [6] provide standardised criteria for assessing sustainable urban development at the neighbourhood level [7, 8, 9], (often qualified as Neighborhood Sustainability Assessments - NSA [10] and NSA Tools (NSATs) [11]) which are helpful in evaluating initiatives to achieve certification [12], but from a strategic perspective, these frameworks often focus primarily on specific (social or environmental) aspects and may overlook broader social and economic dimensions of sustainability
@@ 2.2 ISO 37101: Structure, purpose, and applications
The series containing ISO 37101 [19] , which was introduced in 2016, represents a more holistic approach to assessing urban sustainability, including mapping to a set of urban indicators [20], specified for example in ISO37120 and operationalized in cities indicators datasets [21, 22, 23]. Its structure is based on six purposes (attractiveness, preservation and improvement of the environment, resilience, responsible resource use, social cohesion, and well-being) and twelve issues (from governance to biodiversity). provides a comprehensive matrix for assessing urban initiatives, as well as specific indicators [24]. This approach was recently explored by the French public expertise agency for the ecological transition and regional planning (CEREMA) [25, 26]. Thanks to the flexibility of the standard, it can be adapted to different scales, from individual projects to city-wide strategies. However, the relative complexity of the framework can make widespread adoption difficult, especially in resource-constrained urban environments
@@ 3 Methodology
We selected gpt-3.5-turbo as the primary LLM, with additional tests using other local models (llama2, mistral, …). The choice of gpt-3.5-turbo was based on its proven ability to understand complex relationships and generate nuanced responses, as well as its ability to use “functions calling” to obtain JSON-structured data. This approach can later be ported to more recent models (in particular the latest OpenAI gpt-4o-mini model). Its large knowledge base, trained on various datasets, makes it well suited for interpreting different urban project descriptions from a general perspective. We also considered factors such as the accessibility of the API, the consistency of the results and the possibilities for fine-tuning. We tried locally-runned models as secondary models to allow to compare the results and qualitatively assess the generalisability of our approach across different LLMs, but rapidly discarded this because of the speed and usability across large datasets.
A critical component of this LLM-based approach the development of a custom prompt that summarises the essence of the ISO 37101 framework. This prompt was iteratively refined to guide the LLM in the use of the 6x12 framework, including a reworked, custom set of definitions for the 18 items. The prompt includes concise definitions of the six sustainability purposes and twelve issues, instructions for identifying relevant aspects of the input text, guidelines for assessing the project’s contribution to each purpose-issue intersection, and a structured format for the output of the assessment results, qualified in the functions calling part of the request. Prompts have been reworked through testing with a variety of urban project descriptions to ensure consistency and accuracy of LLM interpretations.
@@
  • Creation of visualisations and user-friendly reports in Excel format.
4 Results
@@ 4.1.2 Qualitative analysis and control of LLM outputs
Figure 1: A review from an urban project in Grenoble, by Ecocités, showing alignment of tenderers to the needs of the city [31].
@@ 4.2.1 Paris
Moreover, based on this dataset, we developed a catalogue of "sustainable ideas" in which we have categorised exemplary projects under each purpose-issue overlap [33]. This can catalogue serve as a resource for urban planners and citizens and provides inspiration for future participatory budgeting proposals focused on sustainability goals.
@@ 4.2.2 The PROBONO project
Figure 2: Review of the PROBONO [30] activities.
@@ 5.2 Comparison with traditional human-driven assessment methods
The advantages of an LLM-based assistant versus fully human reviews are twofold. The first is the significant time saving compared to traditional manual assessments (a few seconds compared to around 30 minutes per project for an overview of the project). This efficiency gain opens up opportunities for more comprehensive and frequent sustainability assessments, allowing communities to assess a much wider range of initiatives and update their assessments more regularly. The second aspect is the consistency of automated assessments, which offsets another limitation of manual methods, namely the potential variability between assessors. This approach provides a standardised interpretation that can be particularly useful for comparing projects in different contexts or time periods. LLM have also enabled the use of virtual citizens [34], with obvious limitations. Nevertheless, the use of AI in urban planning decision-making processes raises important ethical considerations [35]. It is essential to ensure that LLMs complement, rather than replace, human expertise. Furthermore, while LLMs are excellent at recognising explicit relationships, they occasionally overlook subtle, context-specific factors that can be easily identified by human experts. There is therefore a risk that important qualitative factors or local knowledge that cannot be captured in standardised frameworks will be marginalised.
@@ 5.5 Applicability to solutions identification
This exchange of ideas and solutions can be powerful as it is based on the overview understanding of city services and sustainability impacts rather than superficial similarities. For example, a community’s need for greater social cohesion around environmental preservation may lead to the discovery of innovative green space initiatives that have fostered community engagement elsewhere, but also identify already existing projects, for which the ’other’ impacts were not previously considered. This approach can accelerate the learning process for communities stakeholder, and the adaptation and scaling of proven solutions in different urban contexts. It also encourages a collaborative and globally focssed approach to urban problem solving, where communities can learn from and build on the experiences of others.
@@ 7 Conclusion
Furthermore, the adaptability of our approach in different contexts - from open citizen data in the Paris participatory budget to specific project information in the context of PROBONO activities - demonstrates its versatility and potential for wide-ranging applications in urban planning and governance. This flexibility suggests that the method could be useful for cities of different sizes, cultural contexts and stages of development. It can offer for example a consistent way to present the alignment of different certification schemes mentioned earlier to the ambitions of the community, allowing the decision makers to opt for the certification scheme that best aligns with their priorities.